# Robust Bayesian Inference for Berkson and Classical Measurement Error Models

### v2

@@ B.1 Proofs of theoretical results for the Berkson error model
Lemma B.4.
(Berkson) Let $\mathbb{P}^{0,n}:=\operatorname{\mathbb{P}}_{X}^{n}\operatorname{\mathbb{P}}_{Y}^{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{(x_{i},y_{i})}$ and $\mathbb{P}_{XY}=\operatorname{\mathbb{P}}_{X}\operatorname{\mathbb{P}}_{Y}^{n}=\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{\infty}\xi_{j}^{i}\delta_{(z_{j},y_{i})}$ where $z^{i}_{1:\infty}\overset{iid}{\sim}\operatorname{\mathbb{F}}^{\prime}_{w_{i}},\xi^{i}_{1:\infty}\sim\operatorname{\operatorname{GEM}}(c+1)$. Then under assumptions 1-3:
@@
$\quad\quad\quad\leq\frac{1}{\sqrt{n}}+\frac{1}{\sqrt{c+2}}+\frac{c}{c+1}\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\operatorname{\mathbb{F}}_{N})+\frac{1}{c+1}\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})$
@@
Since $\operatorname{\mathbb{P}}_{Y}^{n}$ does not depend on $X$, it admits the trivial conditional mean embedding operator $\mathcal{C}_{Y}:\operatorname{\mathcal{H}_{k_{X}}}\rightarrow\operatorname{\mathcal{H}_{k_{Y}}}$ defined by $\mathcal{C}_{Y}f=\mu_{\operatorname{\mathbb{P}}^{n}_{Y}}$ for any $f\in\operatorname{\mathcal{H}_{k_{X}}}$. Hence, for any $f\in\operatorname{\mathcal{H}_{k_{X}}}$, and using Assumption 4.0.2, we have:
$\sup_{y_{1:n}}\|\mu_{\operatorname{\mathbb{P}}^{n}_{Y}}\|^{2}_{\operatorname{\mathcal{H}_{k_{Y}}}}=\sup_{y_{1:n}}\left\|\frac{1}{n}\sum_{i=1}^{n}k_{Y}(y_{i},\cdot)\right\|^{2}_{\operatorname{\mathcal{H}_{k_{Y}}}}=\sup_{y_{1:n}}\left(\frac{1}{n^{2}}\sum_{i=1}^{n}k_{Y}(y_{i},y_{i})+\frac{2}{n^{2}}\sum_{i\neq j}k_{Y}(y_{i},y_{j})\right)=1$
hence $\sup_{y_{1:n}}\|\mu_{\operatorname{\mathbb{P}}^{n}_{Y}}\|_{\operatorname{\mathcal{H}_{k_{Y}}}}\leq 1$ and $\|\mathcal{C}_{Y}\|_{o}\leq 1$. Therefore,
$\mathbb{E}_{(x,y)_{1:n}\sim\operatorname{\mathbb{P}}^{0}}[\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}_{k}(\mathbb{P}^{0,n},\mathbb{P}_{XY})]]$ $\leq\|\mathcal{C}_{Y}\|_{o}\mathbb{E}_{x_{1:n}\sim\operatorname{\mathbb{P}}_{X}^{0,n}}[\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{P}}_{X}^{n},\operatorname{\mathbb{P}}_{X})]]$
$\leq\mathbb{E}_{x_{1:n}\sim\operatorname{\mathbb{P}}_{X}^{0,n}}[\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{P}}_{X}^{n},\operatorname{\mathbb{P}}_{X})]]$
$\leq\mathbb{E}_{x_{1:n}\sim\operatorname{\mathbb{P}}_{X}^{0,n}}[\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{P}}_{X}^{0,n},\operatorname{\mathbb{P}}_{X})+\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{P}}_{X}^{0,n},\operatorname{\mathbb{P}}_{X}^{n})]]$
$\leq\mathbb{E}_{x_{1:n}\sim\operatorname{\mathbb{P}}_{X}^{0,n}}[\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{P}}_{X}^{0,n},\operatorname{\mathbb{P}}_{X})]]+\frac{1}{\sqrt{n}}$
$\leq\frac{1}{\sqrt{n}}+\frac{1}{\sqrt{c+2}}+\frac{c}{c+1}\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\operatorname{\mathbb{F}}_{N})$
$\quad\quad+\frac{1}{c+1}\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})$
where we used Lemma B.3 for the first inequality, the triangle inequality for the third inequality and Lemma B.2 with $k_{X}^{2}$ for the last inequality. The fact that $k_{X}^{2}$ satisfies the required conditions of Lemma B.2 follows from Assumptions 4.0.2 and 4.0.3. In particular, if $|k_{X}|\leq 1$ then $|k^{2}_{X}|\leq 1$ and if $k_{X}$ is translation invariant then $k^{2}_{X}(x,x^{\prime})=\psi(x-x^{\prime})\psi(x-x^{\prime})\forall x,x^{\prime}\in\mathcal{X}$ and $k^{2}_{X}$ is also translation invariant. Finally, for the fourth inequality, we used Lemma B.1 with $S_{i}=(x_{i})\sim Q_{i}:=\operatorname{\mathbb{P}}^{0}_{X\>|\>w_{i}}$. ∎
Lemma B.5.
(Berkson) Let $\operatorname{\mathbb{P}}^{0}_{\theta}:=\operatorname{\mathbb{P}}_{X}^{0,n}\operatorname{\mathbb{P}}_{g(\theta,\cdot)}$ and $\operatorname{\mathbb{P}}_{\theta}:=\operatorname{\mathbb{P}}_{X}\operatorname{\mathbb{P}}_{g(\theta,\cdot)}$. Under Assumptions 1-4, for any $\theta\in\Theta$:
$\mathbb{E}_{x_{1:n}\sim\operatorname{\mathbb{P}}_{X}^{0,n}}[\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{X}^{0,n}\operatorname{\mathbb{P}}_{g(\theta,\cdot)},\operatorname{\mathbb{P}}_{X}\operatorname{\mathbb{P}}_{g(\theta,\cdot)})]]$
@@
$\quad\quad\quad\leq\Lambda\left(\frac{1}{\sqrt{c+2}}+\frac{c}{c+1}\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\operatorname{\mathbb{F}}_{N})+\frac{1}{c+1}\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})\right)$
@@
Using Lemma B.3 and Assumption 4 we have that for any $\theta\in\Theta$:
$\mathbb{E}_{x_{1:n}\sim\operatorname{\mathbb{P}}_{X}^{0,n}}[\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{X}^{0,n}\operatorname{\mathbb{P}}_{g(\theta,\cdot)}),\operatorname{\mathbb{P}}_{X}\operatorname{\mathbb{P}}_{g(\theta,\cdot)}]]$ $\leq\Lambda\mathbb{E}_{x_{1:n}\sim\operatorname{\mathbb{P}}_{X}^{0,n}}[\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{P}}_{X}^{0,n},\operatorname{\mathbb{P}}_{X})]]$
and the result follows by Lemma B.2 for $k^{2}_{X}$. ∎

### v3

@@ D.1 Proofs of theoretical results for the Berkson error model
Lemma D.4.
(Berkson) Let $\mathbb{P}^{0,n}:=\operatorname{\mathbb{P}}_{X}^{n}\times\operatorname{\mathbb{P}}_{Y}^{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{(x_{i},y_{i})}$ and $\mathbb{P}_{XY}=\operatorname{\mathbb{P}}_{X}\times\operatorname{\mathbb{P}}_{Y}^{n}=\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{\infty}\xi_{j}^{i}\delta_{(z_{j},y_{i})}$ where $z^{i}_{1:\infty}\overset{iid}{\sim}\operatorname{\mathbb{F}}^{\prime}_{w_{i}},\xi^{i}_{1:\infty}\sim\operatorname{\operatorname{GEM}}(c+1)$. Then under Assumptions 4.0.1-4.0.3:
@@
$\leq\frac{1}{\sqrt{n(c+2)}}+\frac{c}{c+1}\sqrt{{\text{Var}}_{\mathcal{H}_{k_{X}}}[\operatorname{\mathbb{F}}_{N}^{0}]+\operatorname{\operatorname{MMD}}^{2}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\operatorname{\mathbb{F}}_{N})}+\frac{\sqrt{2}}{c+1}\sqrt{\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})}.$
@@
By the triangle inequality, we have:
@@
$\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0,n},\operatorname{\mathbb{P}}_{XY})$ $\leq\operatorname{\operatorname{MMD}}_{k}\left(\operatorname{\mathbb{P}}^{0,n},\frac{1}{n}\sum_{i=1}^{n}\mathbb{F}^{\prime}_{w_{i}}\times\delta_{y_{i}}\right)+\operatorname{\operatorname{MMD}}_{k}\left(\operatorname{\mathbb{P}}_{X,Y},\frac{1}{n}\sum_{i=1}^{n}\mathbb{F}^{\prime}_{w_{i}}\times\delta_{y_{i}}\right).$ (32)
Let $\phi_{X},\phi_{Y}$ be the feature maps associated with $k_{X}$ and $k_{Y}$ respectively. For the first term, we have:
$\operatorname{\operatorname{MMD}}_{k}\left(\operatorname{\mathbb{P}}^{0,n},\frac{1}{n}\sum_{i=1}^{n}\mathbb{F}^{\prime}_{w_{i}}\times\delta_{y_{i}}\right)$
$=\left\|\frac{1}{n}\sum_{i}\left[\phi_{X}(x_{i})\otimes\phi_{Y}(y_{i})-\frac{c}{c+1}\mu_{\mathbb{F}_{w_{i}}}\otimes\phi_{Y}(y_{i})-\frac{1}{c+1}\phi_{X}(w_{i})\otimes\phi_{Y}(y_{i})\right]\right\|_{\mathcal{H}_{k}}$
$=\left\|\frac{1}{n}\sum_{i}\left[\frac{c}{c+1}(\phi_{X}(x_{i})-\mu_{\mathbb{F}_{w_{i}}})\otimes\phi_{Y}(y_{i})+\frac{1}{c+1}(\phi_{X}(x_{i})-\phi_{X}(w_{i}))\otimes\phi_{Y}(y_{i})\right]\right\|_{\mathcal{H}_{k}}$
$\leq\frac{1}{n}\sum_{i}\left[\frac{c}{c+1}\left\|(\phi_{X}(x_{i})-\mu_{\mathbb{F}_{w_{i}}})\otimes\phi_{Y}(y_{i})\right\|_{\mathcal{H}_{k}}\right$.
$\left.\quad+\frac{1}{c+1}\left\|(\phi_{X}(x_{i})-\phi_{X}(w_{i}))\otimes\phi_{Y}(y_{i})\right\|_{\mathcal{H}_{k}}\right]$
$=\frac{1}{n}\sum_{i}\left[\frac{c}{c+1}\left\|(\phi_{X}(x_{i})-\mu_{\mathbb{F}_{w_{i}}})\right\|_{\mathcal{H}_{k_{X}}}\left\|\phi_{Y}(y_{i})\right\|_{\mathcal{H}_{k_{Y}}}\right]$
$\quad+\frac{1}{n}\sum_{i=1}^{n}\left[\frac{1}{c+1}\left\|(\phi_{X}(x_{i})-\phi_{X}(w_{i}))\right\|_{\mathcal{H}_{k_{X}}}\left\|\phi_{Y}(y_{i})\right\|_{\mathcal{H}_{k_{Y}}}\right].$
Now by taking expectations and using the Cauchy-Schwarz inequality, we get:
$\mathbb{E}_{w_{i},x_{i},y_{i}}\left[\operatorname{\operatorname{MMD}}_{k}\left(\operatorname{\mathbb{P}}^{0,n},\frac{1}{n}\sum_{i=1}^{n}\mathbb{F}^{\prime}_{w_{i}}\times\delta_{y_{i}}\right)\right]$
$\leq\frac{c}{n(c+1)}\sum_{i=1}^{n}\mathbb{E}_{w_{i},x_{i},y_{i}}[\|\phi_{X}(x_{i})-\mu_{\operatorname{\mathbb{F}}_{w_{i}}}\|_{\mathcal{H}_{k_{X}}}\|\phi_{Y}(y_{i})\|_{\mathcal{H}_{k_{Y}}}]$
$\quad+\frac{1}{n(c+1)}\sum_{i=1}^{n}\mathbb{E}_{w_{i},x_{i},y_{i}}[\left\|(\phi_{X}(x_{i})-\phi_{X}(w_{i}))\right\|_{\mathcal{H}_{k_{X}}}\left\|\phi_{Y}(y_{i})\right\|_{\mathcal{H}_{k_{Y}}}]$
$\leq\frac{c}{n(c+1)}\sum_{i}\sqrt{\mathbb{E}_{w_{i},x_{i}}[\|\phi_{X}(x_{i})-\mu_{F_{w_{i}}}\|_{\mathcal{H}_{k_{X}}}^{2}]}\sqrt{\mathbb{E}_{w_{i},x_{i},y_{i}}[\|\phi_{Y}(y_{i})\|_{\mathcal{H}_{k_{Y}}}^{2}]}$
$\quad+\frac{1}{n(c+1)}\sum_{i}\sqrt{\mathbb{E}_{w_{i},x_{i}}[\|\phi_{X}(x_{i})-\phi_{X}(w_{i})\|_{\mathcal{H}_{k_{X}}}^{2}]}\sqrt{\mathbb{E}_{w_{i},x_{i},y_{i}}[\|\phi_{Y}(y_{i})\|_{\mathcal{H}_{k_{Y}}}^{2}]}$
$\leq\frac{c}{n(c+1)}\sum_{i}\sqrt{\mathbb{E}_{w_{i},x_{i}}[\|\phi_{X}(x_{i})-\mu_{F_{w_{i}}}\|_{\mathcal{H}_{k_{X}}}^{2}]}$
$\quad+\frac{1}{n(c+1)}\sum_{i}\sqrt{\mathbb{E}_{w_{i},x_{i}}[\|\phi_{X}(x_{i})-\phi_{X}(w_{i})\|_{\mathcal{H}_{k_{X}}}^{2}]}.$ (33)
For the second term, by Assumption 4.0.3 we have that $k_{X}$ is translation invariant so $\exists\psi_{X}$ such that $k_{X}(x,x^{\prime})=\psi_{X}(x-x^{\prime})$ hence:
$\mathbb{E}_{w_{i},x_{i}}[\|\phi_{X}(x_{i})-\phi_{X}(w_{i})\|_{\mathcal{H}_{k_{X}}}^{2}]$
$=\operatorname{\mathbb{E}}_{\nu\sim\operatorname{\mathbb{F}}^{0}_{N}}[\|\phi_{X}(w_{i}+\nu)-\phi_{X}(w_{i})\|^{2}_{\mathcal{H}_{k_{X}}}]$
$=\operatorname{\mathbb{E}}_{\nu\sim\operatorname{\mathbb{F}}^{0}_{N}}[k_{X}(w_{i}+\nu,w_{i}+\nu)+k_{X}(w_{i},w_{i})-2k_{X}(w_{i}+\nu,w_{i})]$
$=2\psi_{X}(0)-2\operatorname{\mathbb{E}}_{\nu\sim\operatorname{\mathbb{F}}_{N}^{0}}[\psi_{X}(\nu)]$
$=2-2\operatorname{\mathbb{E}}_{\nu\sim\operatorname{\mathbb{F}}_{N}^{0}}[\psi_{X}(\nu)]$
where in the last equality we used the fact that $\psi_{X}(0)=\sup_{x,x^{\prime}\in\mathcal{X}}k_{X}(x,x^{\prime})=1$ since $k_{X}$ is positive definite and hence $\psi_{X}$ needs to have its supremum at 0. We hence also have that:
$\operatorname{\operatorname{MMD}}_{k_{X}}^{2}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})$ $=\|\operatorname{\mathbb{E}}_{\nu\sim\operatorname{\mathbb{F}}_{N}^{0}}[\phi_{X}(\nu)]-\phi_{X}(0)\|^{2}_{\mathcal{H}_{k_{X}}}$
$=\|\operatorname{\mathbb{E}}_{\nu\sim\operatorname{\mathbb{F}}_{N}^{0}}[\phi_{X}(\nu)]\|^{2}_{\mathcal{H}_{k_{X}}}-2\operatorname{\mathbb{E}}_{\nu\sim\operatorname{\mathbb{F}}_{N}^{0}}[\psi_{X}(\nu)]+\psi_{X}(0)$
$=(1-\operatorname{\mathbb{E}}_{\operatorname{\mathbb{F}}_{N}^{0}}[\phi_{X}(N)])^{2}$
$=\frac{1}{4}(2-2\operatorname{\mathbb{E}}_{\nu\sim\operatorname{\mathbb{F}}_{N}^{0}}[\psi_{X}(\nu)])^{2}$
$=\frac{1}{4}\left(\mathbb{E}_{w_{i},x_{i}}[\|\phi_{X}(x_{i})-\phi_{X}(w_{i})\|_{\mathcal{H}_{k_{X}}}^{2}]\right)^{2}$
and hence
$\mathbb{E}_{w_{i},x_{i}}[\|\phi_{X}(x_{i})-\phi_{X}(w_{i})\|_{\mathcal{H}_{k_{X}}}^{2}]=2\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0}).$ (34)
For the first term, note that
$\operatorname{\mathbb{E}}_{w_{i},x_{i}}[\|\phi_{X}(x_{i})-\mathbb{E}_{x\sim\operatorname{\mathbb{F}}_{w_{i}}}[\phi_{X}(x)\|^{2}]$
$=\operatorname{\mathbb{E}}_{x_{i}\sim\operatorname{\mathbb{P}}^{0}_{X|w_{i}}}[k(x_{i},x_{i})]-2\operatorname{\mathbb{E}}_{x_{i}\sim\operatorname{\mathbb{P}}^{0}_{X|w_{i}}}[\left<\phi_{X}(x_{i}),\operatorname{\mathbb{E}}_{x\sim\operatorname{\mathbb{F}}_{w_{i}}}\phi_{X}(x)\right$>]
$\quad\quad+\left<\operatorname{\mathbb{E}}_{\operatorname{\mathbb{F}}_{w_{i}}}[\phi_{X}(x)],\operatorname{\mathbb{E}}_{\operatorname{\mathbb{F}}_{w_{i}}}[\phi_{X}(x)]\right>_{\mathcal{H}_{k_{X}}}$
$=\operatorname{\mathbb{E}}_{\nu\sim\operatorname{\mathbb{F}}_{N}^{0}}[k_{i}(\nu,\nu)]-2\operatorname{\mathbb{E}}_{\nu\sim\operatorname{\mathbb{F}}_{N}^{0}}[\left<\phi_{i}(\nu),\operatorname{\mathbb{E}}_{\nu^{\prime}\sim\operatorname{\mathbb{F}}_{N}}\phi_{i}(\nu^{\prime})\right>_{\mathcal{H}_{k_{X}}}]$
$\quad\quad+\left<\operatorname{\mathbb{E}}_{\operatorname{\mathbb{F}}_{N}}[\phi_{i}(\nu)],\operatorname{\mathbb{E}}_{\operatorname{\mathbb{F}}_{N}}[\phi_{i}(\nu)]\right$>
$=\operatorname{\mathbb{E}}_{\nu\sim\operatorname{\mathbb{F}}_{N}^{0}}[k_{i}(\nu,\nu)]-\left<\mu_{\operatorname{\mathbb{F}}_{N}^{0}},\mu_{\operatorname{\mathbb{F}}_{N}^{0}}\right>+\operatorname{\operatorname{MMD}}^{2}_{k_{i}}(\operatorname{\mathbb{F}}_{N}^{0},\operatorname{\mathbb{F}}_{N})$
$={\text{Var}}_{\mathcal{H}_{k_{i}}}[\operatorname{\mathbb{F}}_{N}^{0}]+\operatorname{\operatorname{MMD}}^{2}_{k_{i}}(\operatorname{\mathbb{F}}_{N}^{0},\operatorname{\mathbb{F}}_{N})$
$={\text{Var}}_{\mathcal{H}_{k_{X}}}[\operatorname{\mathbb{F}}_{N}^{0}]+\operatorname{\operatorname{MMD}}^{2}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\operatorname{\mathbb{F}}_{N})$
where $\phi_{i}$ denotes the feature map of $k_{i}$ which is defined as $k_{i}(\nu,\nu^{\prime}):=k_{X}(w_{i}+\nu,w_{i}+\nu^{\prime})$ and the last equality follows from the fact that $k_{X}$ is translation invariant. We further used the notation: ${\text{Var}}_{\mathcal{H}_{k_{X}}}[\operatorname{\mathbb{F}}_{N}^{0}]:=\operatorname{\mathbb{E}}_{\nu\sim\operatorname{\mathbb{F}}_{N}^{0}}[k_{X}(\nu,\nu)]-\left<\mu_{\operatorname{\mathbb{F}}_{N}^{0}},\mu_{\operatorname{\mathbb{F}}_{N}^{0}}\right>_{\mathcal{H}_{k_{X}}}$ which is the effective RKHS variance of $\operatorname{\mathbb{F}}_{N}^{0}$. When there is no ME, this term will go to zero, whereas otherwise it will quantify the error induced by higher ME distribution variances, making the problem harder. Substituting into (33) we obtain:
$\mathbb{E}_{(x,y)_{1:n}\sim\operatorname{\mathbb{P}}^{0}}\left[\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}\left[\operatorname{\operatorname{MMD}}_{k}\left(\operatorname{\mathbb{P}}^{0,n},\frac{1}{n}\sum_{i=1}^{n}\mathbb{F}^{\prime}_{w_{i}}\times\delta_{y_{i}}\right)\right]\right]$
$\leq\frac{c}{c+1}\sqrt{{\text{Var}}_{\mathcal{H}_{k_{X}}}[\operatorname{\mathbb{F}}_{N}^{0}]+\operatorname{\operatorname{MMD}}^{2}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\operatorname{\mathbb{F}}_{N})}+\frac{\sqrt{2}}{c+1}\sqrt{\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})}.$
We now turn to the second term of (32). Expanding the expected MMD we obtain:
$\mathbb{E}_{(x,y)_{1:n}\sim\operatorname{\mathbb{P}}^{0}}\left[\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}[\left[\operatorname{\operatorname{MMD}}_{k}^{2}(\operatorname{\mathbb{P}}_{XY},\frac{1}{n}\sum_{i=1}^{n}\mathbb{F}^{\prime}_{w_{i}}\times\delta_{y_{i}})\right]\right]$
$=\mathbb{E}_{(x,y)_{1:n}\sim\operatorname{\mathbb{P}}^{0}}\left[\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}\left[\left\|\frac{1}{n}\sum_{i=1}^{n}\left(\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\otimes\phi_{Y}(y_{i})-\mu_{\operatorname{\mathbb{P}}^{i}}\otimes\phi_{Y}(y_{i})\right)\right\|^{2}\right]\right]$
$=\frac{1}{n^{2}}\sum_{i=1}^{n}\mathbb{E}_{w_{i},x_{i},y_{i}}\mathbb{E}_{\operatorname{\mathbb{P}}^{i}}\left[\|\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\otimes\phi_{Y}(y_{i})-\mu_{\operatorname{\mathbb{P}}^{i}}\otimes\phi_{Y}(y_{i})\|^{2}\right]$
[…]
The cross terms vanish since $\mathbb{E}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\mathbb{P}}^{i}]=\operatorname{\mathbb{F}}^{\prime}_{w_{i}}$ for each $i\in\{1,\dots,n\}$ and conditionally on $(w_{i},w_{j},x_{i},x_{j},y_{i},y_{j})$ $\mu_{\operatorname{\mathbb{P}}^{i}}\otimes\phi_{Y}(y_{i})$ and $\mu_{\operatorname{\mathbb{P}}^{j}}\otimes\phi_{Y}(y_{j})$ are independent and hence:
[…]
$=0.$
For the diagonal terms, we have:
$\mathbb{E}_{w_{i},x_{i},y_{i}}\mathbb{E}_{\operatorname{\mathbb{P}}_{i}}\left[\|\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\otimes\phi_{Y}(y_{i})-\mu_{\operatorname{\mathbb{P}}_{i}}\otimes\phi_{Y}(y_{i})\|^{2}\right]$
$=\mathbb{E}_{w_{i},x_{i},y_{i}}\mathbb{E}_{\operatorname{\mathbb{P}}_{i}}\left[\left\|\sum_{j=1}^{\infty}\xi_{j}\left[\phi_{X}(z_{j})\otimes\phi_{Y}(y_{i})-\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}\otimes\phi_{Y}(y_{i})\right]\right\|^{2}\right]$
$=\mathbb{E}_{w_{i},x_{i},y_{i}}\mathbb{E}_{\operatorname{\mathbb{P}}_{i}}\left[\sum_{j=1}^{\infty}\xi_{j}^{2}\|(\phi_{X}(z_{j})-\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}})\otimes\phi_{Y}(y_{i})\|^{2}\right$.
[…]
Let $\Delta_{j}:=(\phi_{X}(z_{j})-\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}})\otimes\phi_{Y}(y_{i})$. Then
$\mathbb{E}_{w_{i},x_{i},y_{i}}\mathbb{E}_{\operatorname{\mathbb{P}}_{i}}[\Delta_{j}]$ $=\mathbb{E}_{w_{i},x_{i},y_{i}}\left[\mathbb{E}_{z_{j}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[(\phi_{X}(z_{j})-\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}})\otimes\phi_{Y}(y_{i})]\right]$
$=\mathbb{E}_{w_{i},x_{i},y_{i}}\left[(\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}-\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}})\otimes\phi_{Y}(y_{i})\right]$
$=0$
and $\Delta_{j}$ is independent of $\Delta_{t}$ conditionally on $w_{i},x_{i},y_{i}$ for $j\neq t$ so the cross terms vanish. Finally, we have:
$\mathbb{E}_{w_{i},x_{i},y_{i}}\mathbb{E}_{\operatorname{\mathbb{P}}_{i}}\left[\|\Delta_{j}\|^{2}\right]$
[…]
$\leq\mathbb{E}_{w_{i},x_{i},y_{i}}\mathbb{E}_{z_{j}\sim\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}[k((z_{j},y_{i}),(z_{j},y_{i}))]$
$\leq 1.$
Therefore, putting it all together, we have:
$\mathbb{E}\left[\operatorname{\operatorname{MMD}}_{k}^{2}(\operatorname{\mathbb{P}}_{XY},\frac{1}{n}\sum_{i=1}^{n}\mathbb{F}^{\prime}_{w_{i}}\times\delta_{y_{i}})\right]$
$=\frac{1}{n^{2}}\sum_{i=1}^{n}\mathbb{E}_{w_{i},x_{i},y_{i}}\mathbb{E}_{\operatorname{\mathbb{P}}^{i}}\left[\sum_{j=1}^{\infty}\xi_{j}^{2}\|(\phi_{X}(z_{j})-\mu_{\operatorname{\mathbb{F}}^{\prime}_{w_{i}}})\otimes\phi_{Y}(y_{i})\|^{2}\right]$
$\leq\frac{1}{n^{2}}\sum_{i=1}^{n}\mathbb{E}_{\xi\sim GEM}\left[\sum_{j=1}^{\infty}\xi_{j}^{2}\right]$
$=\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{\infty}\frac{2(c+1)^{j-1}}{(c+3)^{j}(c+2)}$
$=\frac{1}{n(c+2)}.$
By Jensen’s,
$\mathbb{E}\left[\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{XY},\frac{1}{n}\sum_{i=1}^{n}\mathbb{F}^{\prime}_{w_{i}}\times\delta_{y_{i}})\right]\leq\frac{1}{\sqrt{n(c+2)}}$
and the final result is:
$\operatorname{\mathbb{E}}[\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0,n},\operatorname{\mathbb{P}}_{XY})]\leq\frac{1}{\sqrt{n(c+2)}}+\frac{c}{c+1}\sqrt{{\text{Var}}_{\mathcal{H}_{k_{X}}}[\operatorname{\mathbb{F}}_{N}^{0}]+\operatorname{\operatorname{MMD}}^{2}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\operatorname{\mathbb{F}}_{N})}$
$\quad+\frac{\sqrt{2}}{c+1}\sqrt{\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})}.$
Lemma D.5.
(Berkson) Let $\operatorname{\mathbb{P}}^{0}_{\theta}:=\operatorname{\mathbb{P}}_{X}^{0,n}\times\operatorname{\mathbb{P}}_{g(\theta,\cdot)}$ and $\operatorname{\mathbb{P}}_{\theta}:=\operatorname{\mathbb{P}}_{X}\times\operatorname{\mathbb{P}}_{g(\theta,\cdot)}$. Under Assumptions 4.0.1-4.0.4, for any $\theta\in\Theta$:
$\mathbb{E}_{x_{1:n}\sim\operatorname{\mathbb{P}}_{X}^{0,n}}[\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{X}^{0,n}\times\operatorname{\mathbb{P}}_{g(\theta,\cdot)},\operatorname{\mathbb{P}}_{X}\times\operatorname{\mathbb{P}}_{g(\theta,\cdot)})]]$
$\quad\quad\quad\leq\Lambda\left(\frac{1}{\sqrt{n(c+2)}}+\frac{c}{c+1}\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\operatorname{\mathbb{F}}_{N})+\frac{1}{c+1}\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})\right)$
@@
Using Lemma D.3 and Assumption 4 we have that for any $\theta\in\Theta$:
$\mathbb{E}_{x_{1:n}\sim\operatorname{\mathbb{P}}_{X}^{0,n}}[\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{X}^{0,n}\times\operatorname{\mathbb{P}}_{g(\theta,\cdot)},\operatorname{\mathbb{P}}_{X}\times\operatorname{\mathbb{P}}_{g(\theta,\cdot)})]]$
$\quad\quad\leq\Lambda\mathbb{E}_{x_{1:n}\sim\operatorname{\mathbb{P}}_{X}^{0,n}}[\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{P}}_{X}^{0,n},\operatorname{\mathbb{P}}_{X})]]$
and the result follows by Lemma D.2 for $k^{2}_{X}$. ∎